Identify the ΔHf for each product and reactant and solve using the equation:
Example: Determine ΔH for the following reaction:
CS2 (g) + 3O2 ---------- 2 SO2 (g) + 2 H2O (g)
1
Expert's answer
2020-10-26T14:45:19-0400

CS2 (g) + 3O2 (g) = 2SO2 (g) + CO2 (g)

Data:

CS2 (g): 117.1 kJ/mol

3O2 (g): 0

Reactants total: 117.1 kJ/mol

2SO2 (g): 2x(−296.1) kJ/mol

CO2 (g): −393.5 kJ/mol

Products total: 2x(−296.1) + (−393.5) = -985.7 kJ/mol

Produsts - reactants = -985.7 - 117.1 = -1102.8 kJ/mol